OpenStudy (nikolas):

Solve 3 + sin2x = 1 - 6sin2x graphically over the domain 0 ≤ x < 2π. I'm not too sure about how I should start looking at this... am I able to change it to 2 + 7sin2x = 0 ? If so, what would I do next? Thanks.

OpenStudy (anonymous):

if it really means graphically then i guess you should graph \(7\sin(2x)+2\) and see where it crosses the \(x\) axis

OpenStudy (anonymous):

you want \[7\sin(2x)+2=0\] \[7\sin(2x)=-2\] \[\sin(2x)=-\frac{2}{7}\] take the inverse sine and divide by 2 but just graph if

OpenStudy (nikolas):

Thank you! "Solve graphically..." yes, not sure why I didn't graph it, but thanks for explaining as well.